OpenStudy (toxicsugar22):

A doctor who is treated a heart patient wants to presicribe medication to lowere the patients blood pressure. The body's reaction to this medication is a function of the dose admininistered. If the patient takes x milliters of the medication, his blood pressure should decrease by r=f(x) points where f(x)=3x^2-1/3x^3 1. For what values of x is R=0

OpenStudy (anonymous):

So you are trying to find when R = 0, which is when f(x) = 0, which is when 3x^2-1/3x^3 = 0 Start by factoring, what factors are common in both terms on the left of the equation?

OpenStudy (toxicsugar22):

3x

OpenStudy (toxicsugar22):

i guess

OpenStudy (anonymous):

There's more...

OpenStudy (toxicsugar22):

i d k

OpenStudy (anonymous):

ok, not 3, but yes x, and another x can be factored out too: x^2(3 - 1/3x) = 0 Then each part can equal zero to make the whole thing equal zero because they are being multiplied

OpenStudy (anonymous):

...so solve x^2 = 0 and 3-1/3x = 0 and you have both solutions...

OpenStudy (toxicsugar22):

how do i do that

OpenStudy (toxicsugar22):

can u show me

OpenStudy (anonymous):

What values for x can make x^2 = 0 ?

OpenStudy (toxicsugar22):

o

OpenStudy (anonymous):

right, that's one solution

OpenStudy (toxicsugar22):

now how do i do the second equation

OpenStudy (anonymous):

to solve 3 - 1/3x = 0: 3 - 1/3x = 0 add 1/3x to each side 3 - 1/3x + 1/3x = 0 + 1/3x on the left they cancel by subtraction and on the right, the zero is extra: 3 + 0 = 1/3x...

OpenStudy (anonymous):

...then the zero on the right is extra: 3 = 1/3x Then, since x is being multiplied by 1/3, you do the operation that would "undo" that, which is multiply by 3 (same as divide by 1/3) again on both sides: 3*3 = 3*1/3x 9 = 1x 9 = x Second solution

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

from theat question can u help me with one more

OpenStudy (toxicsugar22):

what values of x make sence for this function

OpenStudy (toxicsugar22):

modulas can u help me

OpenStudy (toxicsugar22):

what values of x make sence for this function

OpenStudy (toxicsugar22):

can u help me